OpenStudy (anonymous):

simplify ^3sqrt-27x^3

OpenStudy (anonymous):



myininaya (myininaya):

\[\sqrt[3]{(-3)^3} \cdot \sqrt[3]{x^3}\]

jimthompson5910 (jim_thompson5910):

\[\LARGE \sqrt[3]{-27x^3}\] \[\LARGE \sqrt[3]{(-3x)^3}\] \[\LARGE -3x\] So \[\LARGE \sqrt[3]{-27x^3}=-3x\]

myininaya (myininaya):

\[-3 \cdot x\]